Circular No. 7154

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999bx IN UGC 11391 A. Friedman and W. Li, University of California at Berkeley, on be-

half of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), reports the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of an apparent supernova in UGC 11391. SN 1999bx was discovered on an unfiltered image taken on Apr. 26.5 UT (mag ∼ 16.5) and confirmed on Apr. 27.5 under poor seeing conditions. The new star is located at α = 19^{h}01^{m}41^{s}.44, δ = +40^{o}44′52′′.3 (equinox 2000.0). The host galaxy, UGC 11391, has a peculiar appearance (two nuclei) and may be an interacting galaxy pair. The supernova is sitting in the middle of the western ridge that connects the two nuclei, and is ∼ 2′′.2 west and 14′′.9 north of the southern nucleus. KAIT images of the same field on 1998 Sept. 22.1 (limiting mag ∼ 19.5) showed nothing at the position of SN 1999bx.

NOVA SAGITTARII 1999 P. Garnavich, Harvard-Smithsonian Center for Astrophysics, reports

that spectra of N Sgr 1999 obtained by E. Howell (University of Puerto Rico) and A. Soderberg (Bates College) with the McDonald Observatory 2.1-m telescope on Apr. 27.5 UT show Hα with a clear P-Cyg profile. The emission has a width of 800 km/s (FWHM) and the deep absorption feature has an equivalent width (EW) of 0.7 nm, blueshifted by 1200 km/s. Hβ is observed in absorption with only a small emission peak. Permitted Fe II absorption lines of multiplet 42 are seen at 490.97, 500.39, and 515.41 nm, and strong Na I absorption is also visible with EW = 0.4 nm. The spectrum is consistent with a nova near or before maximum light.

W. Liller, Viña del Mar, Chile, revises his magnitude given on IAUC 7153 (line 16) to V = 7.95. He adds that his previous photograph of the same region, taken with a 85-mm camera lens (+ Tech Pan film + orange filter) on Apr. 21.24 UT, shows no star brighter than V = 11.5 at the position of the new nova. Magnitude estimates via E. O. Waagen, AAVSO: Apr. 27.333 UT, 7.8 (C. Scovil, Stamford, CT, visual); 27.364, V ∼ 7.18 (R. Zissel, South Hadley, MA, CCD).

SUPERNOVA 1999an IN IC 755 Precise CCD position by K. Sárneczky, L. Kiss, and B. Csák for SN

1999an (cf. IAUC 7124), obtained at Konkoly Observatory’s Piszkéstető station on Mar. 19.962 UT: α = 12^{h}01^{m}10^{s}.57, δ = +14^{o}06′12′′.3 (equinox 2000.0), which is 1′′.8 east and 3′′.1 south of the center of IC 755.

1999 April 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green